# Exploring Nonreciprocal Noise Transfer under Onsager-Casimir Symmetry in Synthetic-Field Optomechanics

### v1

@@ II.4 Internal Fluctuations
where $\mathbf{T}(\omega)\coloneq\left[i\omega\mathbf{I}+\mathbf{M}\right]^{-1}$. Applying a similarity transformation which directly diagonalizes the drift matrix $\mathbf{M}$, the explicit frequency dependence of $\mathbf{T}(\omega)$ matrix is found as
@@
Negative frequency component in PSD definition is obtained by Hermitian conjugate relation of operators, $\hat{\mathcal{O}}^{\dagger}(\omega)=\hat{\mathcal{O}}(-\omega)$.Bowen and Milburn (2015) Diagonal elements of $\mathbf{S}(\omega)$ correspond to internal fluctuations of cavity and mechanical modes, respectively. Combining Eqs. (4) and (5)
@@ II.5 Individual contributions to internal and output fluctuations
II.6 Non-reciprocity measure
To quantify the noise flow among the resonators, we define a loop phase-dependent non-reciprocity measure, $I_{\Delta}(\phi_{\ell})$, in terms of integrated spectral densities as
$F_{j\rightarrow i}(\phi_{\ell},\omega)$ $\coloneq$ $T^{\prime}_{R_{i}R_{j}}(\phi_{\ell},\omega)S^{in}_{R_{j}R_{j}}(\phi_{\ell},\omega),$
@@ II.6 Non-reciprocity measure
$I_{\Delta}(\phi_{\ell})$ $\coloneq$ $\int_{-\infty}^{\infty}\left|\left[F_{j\rightarrow i}(\phi_{\ell},\omega)-F_{i\rightarrow j}(\phi_{\ell},\omega)\right]\right|\,d\omega,$ (7)
where $F_{j\rightarrow i}(\phi_{\ell},\omega)$ and $F_{j\rightarrow i}(\phi_{\ell}+\pi,\omega)$ are the time-reversed counterparts of each other and represent the noise transfer from $R_{j}$ to $R_{i}$ quadrature. Input noise PSD matrix is defined as $\mathbf{S}^{in}(\omega)=\langle\mathbf{\hat{R}}^{in}(-\omega)\mathbf{\hat{R}}^{in^{T}}(\omega)\rangle$. Specifically, for the $\phi_{\ell}=0$ or $\pi$ values, time-reversal symmetry is restored, as will be demonstrated in the results to follow.
@@ II.7 Output power spectral density
For sensing applications, it is necessary to determine the output PSD, which relies on the input-output relationship. To explore the sensitivity, a homodyne measurement in optical mode with a homodyne angle $\theta$ is performed, where a rotated quadrature can be measured depending on the homodyne angle. Namely, if $\theta=0^{\circ}$, the position quadrature is measured, whereas if $\theta=90^{\circ}$, the measurement corresponds to the momentum quadrature. By choosing angles other than $0^{\circ}$ or $90^{\circ}$, it becomes possible to measure a rotated quadrature.Sudhir et al. (2017) Output quadratures of the optical cavity and the mechanical resonators can be expressed in terms of the input noises and corresponding susceptibilities, also called transfer functionsBeckey, Carney, and Marocco (2023)
$\hat{R}_{i}^{out}$ $=$ $\chi_{R_{i}Y_{a}}\hat{Y}_{a}^{in}+\chi_{R_{i}X_{a}}\hat{X}_{a}^{in}+\chi_{R_{i}X_{b_{1}}}\hat{X}_{b_{1}}^{in}+\chi_{R_{i}Y_{b_{1}}}\hat{Y}^{in}_{b_{1}}$
$+\chi_{R_{i}X_{b_{2}}}\hat{X}_{b_{2}}^{in}+\chi_{R_{i}Y_{b_{2}}}\hat{Y}_{b_{2}}^{in}.$
@@
$\mathbf{S}^{out}(\omega)=\langle\mathbf{\hat{R}}^{out}(-\omega)\mathbf{\hat{R}}^{out^{T}}(\omega)\rangle.$
@@
$\left[i\omega\mathbf{I}+\mathbf{M}\right]\mathbf{\hat{R}}(\omega)$ $=$ $-\mathbf{D}\mathbf{\hat{R}}^{in}(\omega).$ (8)
For that purpose, we use a slightly modified version of Eq. (4) where the coefficients contained in $\mathbf{\hat{N}}(\omega)$ are transferred into a separate matrix, for the sake of simplicity in calculations. Afterwards, input-output relationship is applied to Eq. (8), Clerk et al. (2010)
$\mathbf{\hat{R}}^{out}(\omega)$ $=$ $\mathbf{\hat{R}}^{in}(\omega)-\mathbf{D}\mathbf{\hat{R}}(\omega),$
$\mathbf{\hat{R}}^{out}(\omega)$ $=$ $\mathbf{P}(\omega)\mathbf{\hat{R}}^{in},$
@@ III The parameter set and its experimental relevance
In this section, we discuss the experimental aspects of our theoretical work before presenting our parameter set. In practical applications, intercoupling two mechanical modes and thereby realizing synthetic magnetism through a loop phase is an experimental challenge.Fang et al. (2017); Del Pino, Slim, and Verhagen (2022); Xu et al. (2019) In our study, individual coupling phases can be gauged out so that only the global loop phase remains. Sütlüoğlu and Bulutay (2021) Hence, this overall loop phase can be realized by embodying it in the intermechanical coupling constant. For instance, coupling can be conveyed through the optical radiation field whose carrier phase is locked to the pump laser with a controllable phase shift via integrated electro-optic and thermo-optic modulators.Gil-Santos et al. (2017); Zhang et al. (2012)
The parameters used in this article are adopted from a seminal experimental work, which substantiates the realizability of our optomechanical system.Safavi-Naeini et al. (2011) However, we should note that this study contained a single mechanical resonator, thus no intermechanical coupling constant, unlike our work. To reduce the parameter space, we consider identical mechanical resonators with the same losses and resonance frequencies. For this purpose, we set, $g_{1,2}\in\mathbb{R}$ with $g_{1}/(2\pi)=g_{2}/(2\pi)=800$ kHz, $\omega_{m}/(2\pi)=3.75$ GHz, $\kappa/(2\pi)=900$ MHz, equal amount of mechanical losses $\gamma_{1}=\gamma_{2}=5\times 10^{-4}\,\omega_{m}$, optical cavity escape efficiency $\eta=0.5$, laser power $P$ = 0.125 mW with a wavelength of 1550 nm, and detuning $\Delta=\omega_{m}$ (red-detuned pumping). This corresponds to the resolved-sideband regime $\kappa<\omega_{m}$ and weak-coupling limit $g_{1,2}\ll\kappa$. This parameter set possesses two EPs, $|\mu_{EP,1}|\simeq 31.06(\gamma_{1}+\gamma_{2})$ and $|\mu_{EP,2}|\simeq 41.6(\gamma_{1}+\gamma_{2})$, where both eigenvalues and the eigenvectors coalesce in the parameter space. Numerical values for the EPs are found from the root loci of the eigenvalues, $(z\left(|\mu|,\phi_{\ell}\right)=\alpha+i\omega,~\omega>0)$ of the drift matrix $\mathbf{M}$ (for details, see Ref. Sütlüoğlu Ege and Bulutay (2024)) Unless stated otherwise, we choose the intermechanical coupling constant value at the second EP, $\mu_{EP,2}$. Therefore, the optomechanical system operates at the strong coupling regime, i.e., $\mu\gg\gamma_{1,2}$. Okamoto et al. (2013); Luo et al. (2018); Deng et al. (2016); Mathew et al. (2016) We assume the optical cavity mode is in the vacuum state, while both mechanical reservoirs have an equal mean number of phonons, $n_{b_{1}}=n_{b_{2}}=100$; this corresponds to an ambient temperature of T=18.1T=18.1~K.
@@ IV Results
We initially focus on intracavity fluctuations and the backaction noise contribution from the second mechanical resonator to the cavity-output PSD. The first mechanical resonator’s contribution to the cavity-output PSD closely matches that of the second resonator, so it is not shown separately. This individual contribution is denoted as $S^{out}_{2\rightarrow{cav}}(\omega)=|P_{Y_{a}Y_{b_{2}}}(\omega)|^{2}S^{in}_{Y_{b_{2}b_{2}}}$. First, we begin by analyzing the intracavity spectrum for loop phases $\phi_{\ell}=0,~\pi/2,~\pi$, and $3\pi/2$ (Fig. 2(a)). As a matter of fact the spectrum for $\phi_{\ell}=3\pi/2$ is identical to that for $\pi/2$. Here, the main observation is that the intracavity noise spectrum can be switched from lower to upper band by tuning the loop-phase from $\phi_{\ell}=0$ to $\pi$ while overall noise power essentially stays the same. This also applies to the cavity output as seen in Fig. 2(b). The fact that the backaction noise spectrum can be swept in frequency by the loop phase is a ramification of backaction being a correlated noise.Yanay, Sankey, and Clerk (2016); Nielsen et al. (2016); Brunelli, Malz, and Nunnenkamp (2019) The backaction noise at $\phi_{\ell}=\pi/2,~3\pi/2$ display the superpositions of $\phi_{\ell}=0$ and $\pi$.
Figure 2: (a) Optical intracavity PSD, (b) noise contribution of the second mechanical resonator to the cavity-output noise spectrum for loop-phases, $\phi_{\ell}=0,~\pi/2,~\pi$, and $3\pi/2$.
Before investigating the non-reciprocity in mechanical resonators, we first examine the internal noise PSD in Fig. 3. It is evident that by appropriately selecting the loop phase as $\phi_{\ell}=\pi/2$ and $3\pi/2$, one can boost the internal noise of the first and second mechanical resonator, respectively. This corroborates our earlier report that, at these loop phases, one of the mechanical resonators can be selectively cooled,Sütlüoğlu Ege and Bulutay (2024) highlighting the role of the synthetic phase in regulating fluctuations.
@@
Figure 5: Non-reciprocity measure $I_{\Delta}(\phi_{\ell})$ with respect to $\phi_{\ell}$ and mechanical coupling constant normalized to $\mu_{EP,2}$.
@@
In Fig. 4, we plot the individual noise transfers for loop phases, $\phi_{\ell}=\pi/2$ and $3\pi/2$: (a) from the second mechanical resonator to the first $S_{2\rightarrow 1}(\omega)=|T^{\prime}_{Y_{b_{1}}Y_{b_{2}}}(\omega)|^{2}S^{in}_{Y_{b_{2}}Y_{b_{2}}}$, and (b) from the first mechanical resonator to the second $S_{1\rightarrow 2}(\omega)=|T^{\prime}_{Y_{b_{2}}Y_{b_{1}}}(\omega)|^{2}S^{in}_{Y_{b_{1}}Y_{b_{1}}}$, respectively. Here, we see that there is a non-reciprocal noise transfer depending on the loop phase. Whenever this phase is $\pi/2$, noise flow from the second to the first mechanical resonator is dominant, and vice versa for $3\pi/2$. This result is also consistent with earlier finding where we achieved a dominant ground state cooling of second and first mechanical resonators when loop phase is chosen as $\phi_{\ell}=\pi/2$ and $\phi_{\ell}=3\pi/2$, respectively.Sütlüoğlu Ege and Bulutay (2024) These figures illustrate the system’s behavior at the second EP, $\mu_{EP,2}$; we refer to Appendix A for additional plots at different mechanical coupling constants. Subsequently, we integrate these spectral contributions to obtain the non-reciprocal noise measure. Hence, we investigate $I_{\Delta}(\phi_{\ell})$ as a function of the mechanical coupling constant normalized to $\mu_{EP,2}$ and the loop phase in Fig. 5. Naively, one would expect a characteristics that is symmetric about $\pi/2$; however, non-reciprocity measure exhibits a pronounced asymmetry, which is caused by the non-trivial loop phase dependence of three-mode coupling.Sütlüoğlu Ege and Bulutay (2024) $I_{\Delta}(\phi_{\ell})$ increases sharply as soon as the loop-phase $\phi_{\ell}$ is detuned even slightly from 0 or $\pi$, reflecting the key role of the broken time-reversal symmetry. As the mechanical coupling constant is increased, the system gradually regains reciprocity, and the non-reciprocal response diminishes, which questions the effect of the EPs. However, their significance becomes evident in Figs. 6 and 7 where we plot the momentum quadrature PSD of the second mechanical resonator for analyzing the internal fluctuations.
@@
Figure 6: PSD of second mechanical resonator with respect to $\phi_{\ell}$ and $\mu$ at $\omega=\omega_{m}$. Horizontal white-dashed lines mark the first and second EPs, $|\mu_{EP,1}|$ and $|\mu_{EP,2}|$.
@@
Figure 7: PSD of second mechanical resonator as a function of frequency (normalized to $\omega_{m}$) and $\mu$ for $\phi_{\ell}=\pi/2$. Horizontal white-dashed lines mark the first and second EPs, $|\mu_{EP,1}|$ and $|\mu_{EP,2}|$.
@@
Analyzing the PSD of internal fluctuations is complicated due to their intricate dependence on multiple parameters such as $\omega$, $\mu$, and $\phi_{\ell}$. Therefore, Figs. 6 and 7 are to be considered as two complementary figures, first one is at $\omega=\omega_{m}$ as a function of $\phi_{\ell}$ at fixed $\mu$, and the second one is at $\phi_{\ell}=\pi/2$ as a function $\omega$ at fixed $\phi_{\ell}$. The behavior of PSD with varying loop-phase in Fig. 6 agrees with Figs. 3 and 4, which confirms that noise in mechanical resonators can be controlled with loop-phase. According to Fig. 5, the non-reciprocity measure decreases when the system approaches EPs, whereas Fig. 7 suggests that sensitivity increases with the increasing $\mu$. Therefore, devices like isolators require a smaller mechanical coupling constant for a non-reciprocal noise flow, but to achieve a better sensitivity (reduced noise PSD at $\omega_{m}$), a system operating closer to the EP will be much more convenient for sensors. Fig. 7 shows that PSD develops a characteristic double-peak structure as $\mu$ approaches $\mu_{EP,1}$, similar to our earlier study where a transition from single to double transparency peak was observed at the EP.Sütlüoğlu and Bulutay (2021) For smaller values of $\mu$, noise becomes dominant at the resonance frequency $\omega_{m}$.
V Conclusion

### v2

@@ II.4 Internal fluctuations
where $\mathbf{T}(\omega)\coloneq\left[i\omega\mathbf{I}+\mathbf{M}\right]^{-1}$. Applying a similarity transformation that directly diagonalizes the drift matrix $\mathbf{M}$, the explicit frequency dependence of $\mathbf{T}(\omega)$ matrix is found as
@@
Negative frequency component in PSD definition is obtained by Hermitian conjugate relation of operators, $\hat{\mathcal{O}}^{\dagger}(\omega)=\hat{\mathcal{O}}(-\omega)$ [68]. Diagonal elements of $\mathbf{S}(\omega)$ correspond to internal fluctuations of cavity and mechanical modes, respectively. Combining Eqs. (7) and (8),
@@ II.5 Individual contributions to internal and output fluctuations
II.6 Nonreciprocity measure
To quantify the noise flow among the resonators, we define a loop phase-dependent nonreciprocity measure, $I_{\Delta}(\phi_{\ell})$, in terms of integrated spectral densities as
$F_{j\rightarrow i}(\phi_{\ell},\omega)$ $\coloneq\bigl|T^{\prime}_{R_{i}R_{j}}(\phi_{\ell},\omega)\bigr|^{2}S^{\mathrm{in}}_{R_{j}R_{j}}(\phi_{\ell},\omega),$
@@ II.6 Nonreciprocity measure
$I_{\Delta}(\phi_{\ell})$ $\coloneq\left|\int_{-\infty}^{\infty}\!\!\left[F_{j\rightarrow i}(\phi_{\ell},\omega)-F_{i\rightarrow j}(\phi_{\ell},\omega)\right]\,d\omega\right|,$ (10)
where $F_{j\rightarrow i}(\phi_{\ell},\omega)$ and $F_{j\rightarrow i}(\phi_{\ell}+\pi,\omega)$ are the time-reversed counterparts of each other and represent the noise transfer from $R_{j}$ to $R_{i}$ quadrature. Input noise PSD matrix is defined as $\mathbf{S}^{\mathrm{in}}(\omega)=\langle\mathbf{\hat{R}}^{\mathrm{in}}(-\omega)\mathbf{\hat{R}}^{\mathrm{in}^{T}}(\omega)\rangle$. This way we characterize nonreciprocity in physical terms as an imbalance in the noise flow between the quadratures $R_{i}$ and $R_{j}$. Notably, for the $\phi_{\ell}=0$ or $\pi$ values, time-reversal symmetry is restored, as will be demonstrated in the results to follow.
@@ II.7 Output power spectral density
For sensing applications, it is necessary to determine the output PSD, which relies on the input-output relationship. To explore the sensitivity, a homodyne measurement in optical mode with a homodyne angle $\theta$ is performed, where a rotated quadrature can be measured depending on the homodyne angle. Namely, if $\theta=0^{\circ}$, the position quadrature is measured, whereas if $\theta=90^{\circ}$, the measurement corresponds to the momentum quadrature. By choosing angles other than $0^{\circ}$ or $90^{\circ}$, it becomes possible to measure a rotated quadrature [69]. Output quadratures of the optical cavity and the mechanical resonators can be expressed in terms of the input noises and corresponding susceptibilities, also called transfer functions [70]:
$\hat{R}_{i}^{\mathrm{out}}$ $=$ $\chi_{R_{i}Y_{a}}\hat{Y}_{a}^{\mathrm{in}}+\chi_{R_{i}X_{a}}\hat{X}_{a}^{\mathrm{in}}+\chi_{R_{i}X_{b_{1}}}\hat{X}_{b_{1}}^{\mathrm{in}}+\chi_{R_{i}Y_{b_{1}}}\hat{Y}^{\mathrm{in}}_{b_{1}}$
$+\chi_{R_{i}X_{b_{2}}}\hat{X}_{b_{2}}^{\mathrm{in}}+\chi_{R_{i}Y_{b_{2}}}\hat{Y}_{b_{2}}^{\mathrm{in}}.$
@@
$\mathbf{S}^{\mathrm{out}}(\omega)=\langle\mathbf{\hat{R}}^{\mathrm{out}}(-\omega)\mathbf{\hat{R}}^{\mathrm{out}^{T}}(\omega)\rangle.$
@@
$\left[i\omega\mathbf{I}+\mathbf{M}\right]\mathbf{\hat{R}}(\omega)$ $=$ $-\mathbf{D}\mathbf{\hat{R}}^{\mathrm{in}}(\omega).$ (11)
For that purpose, we use a slightly modified version of Eq. (7) where the coefficients contained in $\mathbf{\hat{N}}(\omega)$ are transferred into a separate matrix, for the sake of simplicity in calculations. Afterwards, input-output relationship is applied to Eq. (11) [67],
$\mathbf{\hat{R}}^{\mathrm{out}}(\omega)$ $=$ $\mathbf{\hat{R}}^{\mathrm{in}}(\omega)-\mathbf{D}\mathbf{\hat{R}}(\omega),$
$\mathbf{\hat{R}}^{\mathrm{out}}(\omega)$ $=$ $\mathbf{P}(\omega)\mathbf{\hat{R}}^{\mathrm{in}},$
@@ III The parameter set and its experimental relevance
In this section, we discuss the experimental aspects of our theoretical work before presenting our parameter set. In practical applications, intercoupling two mechanical modes and thereby realizing synthetic magnetism through a loop phase is an experimental challenge [71, 72, 73]. In our study, individual coupling phases can be gauged out so that only the global loop phase remains [10]. Hence, this overall loop phase can be realized by embodying it in the intermechanical coupling constant. For instance, coupling can be conveyed through the optical radiation field whose carrier phase is locked to the pump laser with a controllable phase shift via integrated electro-optic and thermo-optic modulators [74, 75].
The parameters used in this article are adopted from a seminal experimental work, which substantiates the realizability of our optomechanical system [76]. However, we should note that this study contained a single mechanical resonator, thus no intermechanical coupling constant, unlike our work. To reduce the parameter space, we consider identical mechanical resonators with the same losses and resonance frequencies. For this purpose, we set, $g_{1,2}\in\mathbb{R}$ with $g_{1}/(2\pi)=g_{2}/(2\pi)=800$ kHz, $\omega_{m}/(2\pi)=3.75$ GHz, $\kappa/(2\pi)=900$ MHz, equal amount of mechanical losses $\gamma_{1}=\gamma_{2}=5\times 10^{-4}\,\omega_{m}$, optical cavity escape efficiency $\eta=0.5$, laser power $P$ = 0.125 mW with a wavelength of 1550 nm, and detuning $\Delta=\omega_{m}$ (red-detuned pumping). This corresponds to the resolved-sideband regime $\kappa<\omega_{m}$ and weak-coupling limit $g_{1,2}\ll\kappa$. This parameter set possesses two EPs, $|\mu_{\mathrm{EP},1}|\simeq 31.06\,(\gamma_{1}+\gamma_{2})$ and $|\mu_{\mathrm{EP},2}|\simeq 41.6\,(\gamma_{1}+\gamma_{2})$, where both eigenvalues and the eigenvectors coalesce in the parameter space. Numerical values for the EPs are found from the root loci of the eigenvalues, $(z\left(|\mu|,\phi_{\ell}\right)=\alpha+i\omega,~\omega>0)$ of the drift matrix $\mathbf{M}$  [16]. Unless stated otherwise, we choose the intermechanical coupling constant value at the second EP, $\mu_{\mathrm{EP},2}$. Therefore, the optomechanical system operates at the strong coupling regime, i.e., $\mu\gg\gamma_{1,2}$ [77, 78, 79, 80]. We assume the optical cavity mode is in the vacuum state, while both mechanical reservoirs have an equal mean number of phonons, $n_{b_{1}}=n_{b_{2}}=100$; this corresponds to an ambient temperature of T=18.1T=18.1~K.
@@ IV Results
We initially focus on intracavity fluctuations and the backaction noise contribution from the second mechanical resonator to the cavity-output PSD. The first mechanical resonator’s contribution to the cavity-output PSD closely matches that of the second resonator, so it is not shown separately. This individual contribution is denoted as $S^{out}_{2\rightarrow{cav}}(\omega)=|P_{Y_{a}Y_{b_{2}}}(\omega)|^{2}S^{in}_{Y_{b_{2}b_{2}}}$. First, we begin by analyzing the intracavity spectrum for loop phases $\phi_{\ell}=0,~\pi/2,~\pi$, and $3\pi/2$ [Fig. 2(a)]. As a matter of fact, the spectrum for $\phi_{\ell}=3\pi/2$ is identical to that for $\pi/2$. Here, the main observation is that the intracavity noise spectrum can be switched from lower to upper band by tuning the loop phase from $\phi_{\ell}=0$ to $\pi$ while overall noise power essentially stays the same. This also applies to the cavity output as seen in Fig. 2(b). The fact that the backaction noise spectrum can be swept in frequency by the loop phase is a ramification of backaction being a correlated noise [81, 82, 83]. The backaction noise at $\phi_{\ell}=\pi/2,~3\pi/2$ displays the superpositions of $\phi_{\ell}=0$ and $\pi$.
Figure 2: (a) Optical intracavity PSD, (b) noise contribution of the second mechanical resonator to the cavity-output noise spectrum for loop phases, $\phi_{\ell}=0,~\pi/2,~\pi$, and $3\pi/2$.
Before investigating the nonreciprocity in mechanical resonators, we first examine the internal noise PSD in Fig. 3. It is evident that by appropriately selecting the loop phase as $\phi_{\ell}=\pi/2$ and $3\pi/2$, one can boost the internal noise of the first and second mechanical resonator, respectively. This corroborates our earlier report that, at these loop phases, one of the mechanical resonators can be selectively cooled,[16] highlighting the role of the synthetic phase in regulating fluctuations.
@@
Figure 5: Nonreciprocity measure $I_{\Delta}(\phi_{\ell})$ as a function of the loop phase $\phi_{\ell}$ (horizontal axis) and mechanical coupling constant normalized to second EP $|\mu/\mu_{EP,2}|$ (vertical axis). Dark-red regions indicate stronger nonreciprocity.
@@
In Fig. 4, we plot the individual noise transfers for loop phases, $\phi_{\ell}=\pi/2$ and $3\pi/2$: (a) from the second mechanical resonator to the first $S_{2\rightarrow 1}(\omega)=|T^{\prime}_{Y_{b_{1}}Y_{b_{2}}}(\omega)|^{2}S^{\mathrm{in}}_{Y_{b_{2}}Y_{b_{2}}}$, and (b) from the first mechanical resonator to the second $S_{1\rightarrow 2}(\omega)=|T^{\prime}_{Y_{b_{2}}Y_{b_{1}}}(\omega)|^{2}S^{\mathrm{in}}_{Y_{b_{1}}Y_{b_{1}}}$, respectively. Here, we see that there is a nonreciprocal noise transfer depending on the loop phase. Whenever this phase is $\pi/2$, noise flow from the second to the first mechanical resonator is dominant, and vice versa for $3\pi/2$. This result is also consistent with earlier finding where we achieved a dominant ground state cooling of second and first mechanical resonators when loop phase is chosen as $\phi_{\ell}=\pi/2$ and $\phi_{\ell}=3\pi/2$, respectively [16]. These figures illustrate the system’s behavior at the second EP, $\mu_{\mathrm{EP},2}$; we refer to Appendix A for additional plots at different mechanical coupling constants. Subsequently, we integrate these spectral contributions to obtain the nonreciprocal noise measure. Hence, we investigate $I_{\Delta}(\phi_{\ell})$ as a function of the mechanical coupling constant normalized to $\mu_{\mathrm{EP},2}$ and the loop phase in Fig. 5. Naively, one would expect a characteristic that is symmetric about $\pi/2$; however, nonreciprocity measure exhibits a pronounced asymmetry, which is caused by the non-trivial loop phase dependence of three-mode coupling; further details are provided in Appendix B. $I_{\Delta}(\phi_{\ell})$ increases sharply as soon as the loop phase $\phi_{\ell}$ is detuned even slightly from 0 or $\pi$, reflecting the key role of the broken time-reversal symmetry. As the mechanical coupling constant is increased, the system gradually regains reciprocity, and the nonreciprocal response diminishes. However, their significance becomes evident in Figs. (6) and (7) where we plot the momentum quadrature PSD of the second mechanical resonator for analyzing the internal fluctuations.
@@
Figure 6: PSD of second mechanical resonator with respect to loop phase $\phi_{\ell}$ (horizontal axis) and mechanical coupling constant normalized to second EP $|\mu/\mu_{\mathrm{EP},2}|$ (vertical axis) at $\omega=\omega_{m}$. Horizontal white-dashed lines mark the first and second EPs, $|\mu_{\mathrm{EP},1}|$ and $|\mu_{\mathrm{EP},2}|$. Dark-red regions correspond to areas of enhanced noise in the PSD.
@@
Figure 7: PSD of second mechanical resonator as a function of frequency (normalized to $\omega_{m}$, horizontal axis) and mechanical coupling constant normalized to second EP $|\mu/\mu_{\mathrm{EP},2}|$ (vertical axis) for $\phi_{\ell}=\pi/2$. Horizontal white-dashed lines mark the first and second EPs, $|\mu_{\mathrm{EP},1}|$ and $|\mu_{\mathrm{EP},2}|$. Dark-red regions correspond to enhanced noise levels, while dark-blue regions indicate reduced noise.
@@
Analyzing the PSD of internal fluctuations is complicated due to their intricate dependence on multiple parameters such as $\omega$, $\mu$, and $\phi_{\ell}$. Therefore, Figs. 6 and 7 are to be considered as two complementary figures, first one is at $\omega=\omega_{m}$ as a function of $\phi_{\ell}$ at fixed $\mu$, and the second one is at $\phi_{\ell}=\pi/2$ as a function $\omega$ at fixed $\phi_{\ell}$. The behavior of PSD with varying loop phase in Fig. 6 agrees with Figs. (3) and (4), which confirms that noise in mechanical resonators can be controlled with loop phase. According to Fig. 5, the nonreciprocity measure decreases when the system approaches EPs, whereas Fig. 7 suggests that sensitivity increases with the increasing $\mu$. Therefore, devices like isolators require a smaller mechanical coupling constant for a nonreciprocal noise flow, but to achieve a better sensitivity (reduced noise PSD at $\omega_{m}$), a system operating closer to the EP will be much more convenient for sensors. Fig. 7 shows that PSD develops a characteristic double-peak structure as $\mu$ approaches $\mu_{EP,1}$, similar to our earlier study where a transition from single to double transparency peak was observed at the EP [10]. For smaller values of $\mu$, noise becomes dominant at the resonance frequency $\omega_{m}$.
V Conclusions